Weather breezy, warm & cloudy.  Starting to rain.  Identified supers ready for extraction next week (porter bee escapes to be added on Thursday).
Hive#1 – Not checked, poss virgin queen
Hive#2 – Not checked, poss virgin queen
Hive#4 – Inspected by Brad.  Brood on 9 frames, question whether sufficient room. 4 supers, 2 ready for extraction.  Swap brood frame for foundation at next inspection.  Queen could do with re-marking. Add super next week.  Some sac brood.
Hive#5 – Inspected by Mel.  Brood + half.  Brood on 1 frame in deep.  Stores and lots of pollen.  Supers not ready for extraction. Super added last week not started to be filled.  Feisty bees.
Hive#6 – not checked.
Hive#7  – inspected by Liz.   Searched for new queen but not seen.  Temper good so poss queenright.  Believe queen in bottom but QE under top deep.  Bottom super ready to extract next week.
Hive #7A – not checked.
Hive #8 – inspected by Richard.  Queen not seen. Brood on 7 frames.  Super not ready for extraction.  Plenty of room.
Hive#15F – inspected by Peter.  QC knocked down.  Queen seen.  Angry bees.  Second super added. Brood on 7 frames.
Hive#15H – inspected by Richard & Mel.  2 X deep (no supers last inspection).  Added super.  Top super ready to extract next week.  Moved a further 3 feet towards Hive 3 site.
Hive 16 B – inspected by Brad.  Very strong Nuc. Brood on 5 frames – no room.  Needs hiving.
Hive 16C – inspected by Richard.  Nuc.  Brood on 3 framnes. Needs super frames replacing with deeps after brood hatched.